OpenStudy (anonymous):

y = f(x) = 8^x Find f(x) when x = 1 ____ 3

OpenStudy (cwrw238):

when x = 1/3 f(x) = 8^(1/3)

OpenStudy (cwrw238):

note 8^(1/3) = cube root of 8

OpenStudy (jdoe0001):

\(\Large {\bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \ 8 \to \square \cdot \square \cdot \square \to \square ^3? }\)

OpenStudy (anonymous):

so its 8 lol

OpenStudy (jdoe0001):

know a number that multiplied 3 times by itself is 8?

OpenStudy (anonymous):

is it 2 the answer

OpenStudy (anonymous):

nooo wait

OpenStudy (jdoe0001):

hmm ok

OpenStudy (anonymous):

its 2

OpenStudy (jdoe0001):

yeap \(\large \bf 8^x\qquad x={\color{brown}{ \frac{1}{3}}}\implies 8^{{\color{brown}{ \frac{1}{3}}}}\implies \sqrt[3]{8^1}\implies \sqrt[3]{2^3}\implies 2 \)

OpenStudy (anonymous):

YESSSSSSSSSSSSSSSSSSSSS'

OpenStudy (anonymous):

thank you to both!

OpenStudy (jdoe0001):

yw